OpenStudy (anonymous):

find the domain, x and y-intercepts, vertical and horizantal asymptotes for the rational function y=(x^3-27)/x^4-2x^3+9x^2-18x

OpenStudy (anonymous):

@jim_thompson5910 Could you help me with this?

jimthompson5910 (jim_thompson5910):

\[\Large y=\frac{x^3-27}{x^4-2x^3+9x^2-18x}\] right?

OpenStudy (aum):

@Mhm120: jim_thompson asked you a question. If you confirm the expression he has written is correct he can continue with the answer.

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

and sorry was the late reply

OpenStudy (aum):

Factor the numerator and the denominator. For the numerator use the identity: \(a^3 - b^3 = (a-b)(a^2 - ab + b^2)\). For the denominator factor out x. Try a few small values of x such as x = 1, -1, 2, -2, etc to see if you can find a root of the remaining expression to see if you can find another factor.

OpenStudy (aum):

\(a^3 - b^3 = (a-b)(a^2 + ab + b^2)\) (plus ab not minus ab)

OpenStudy (anonymous):

sorry, but i don't get the part where you're using the small values to find the remaining expression

OpenStudy (aum):

\( x^4 - 2x^3 + 9x^2 - 18x = x(x^3 - 2x^2+9x-18) \) We want to see if \((x^3 - 2x^2+9x-18) \) can be factored. Try to find if x = -1, +1, -2, +2 will make the expression zero. If x = a makes it zero then (x-a) is a factor.

OpenStudy (aum):

Once we reduce it to a quadratic we have factorization or quadratic formula that we can use to find further real factors, if any. Using the identity we already have a quadratic expression in the numerator.

OpenStudy (anonymous):

oh okay thanks!!

OpenStudy (aum):

You are welcome. \( \Large y=\frac{x^3-27}{x^4-2x^3+9x^2-18x} = \frac{(x-3)(x^2+3x+9)}{x(x-2)(x^2+9)} \) There are no other real factors but there are complex factors which we don't have to worry about. Now you can easily find the domain, x-intercepts, y-intercepts, etc.

OpenStudy (anonymous):

oh okay got it thankyou so much

OpenStudy (aum):

yw.